OpenStudy (anonymous):

Write the solution for the following problem as an ordered pair and two equations: The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Be sure to show all work.

OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x." X + X + Y = 46 (can also be written as 2x + Y = 46) Y + 2X = 46 Y - X = 4 * ----------------- (Need to cancel the Y or the X, but cannot do so, so another step is needed) * (Y - X = 4) / -1 (divide each number by -1) = -1Y + 1X = - 4 Y + 2X = 46 -Y + 1X = -4 ------------------- 3X = 42 (divide both sides by 3) --- --- 3 3 X = 14 (Now solve for Y) y + (2 times 14) = 46 y + 28 = 46 1y = 46-28 1y = 18 (divide both sides by 1) --- --- 1 1 Y= 18 So, 14 + 14 + 18 = 46 x + x + y = 46